OpenStudy (anonymous):

The volume of a rectangular prism is 183 in³. What is the volume of a pyramid which has the same height and a congruent base?

OpenStudy (anonymous):

The equation for finding the volume of a pyramid is: Area of the base * Height * 1/3. The volume of a pyramid with the same base and height is 1/3 the volume of a prism (1/3)183in^3=61in^3.

OpenStudy (anonymous):

your answer is 61 inches

OpenStudy (anonymous):

how you get that (1/3) then times what

OpenStudy (anonymous):

Surface Area = 2Lw + 2wh + 2Lh = 150cm²

OpenStudy (anonymous):

So, if you new sides are trippled

OpenStudy (anonymous):

surface Area = 2(3L)(3w) + 2(3w)(3h) + 2(3L)(3h) = 18Lw + 18wh + 18Lh

OpenStudy (anonymous):

so 18+18+18 = 54

OpenStudy (anonymous):

(1/3)183in^3=61in^3

OpenStudy (anonymous):

volume of pyramid = 1/3 volume of rec prism vol of pyramid = 1/3 * 183 = 61

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

the volume of a pyramid is 1/3 since your rectangular prism is 183 you multiply both of those together and you get 61 inches... might be confusing at first but you get to know after a while

OpenStudy (anonymous):

@rebecca1233 Do u get it?

OpenStudy (anonymous):

yes. thanks

OpenStudy (anonymous):

No problem become a fan of mines if you need help on anything else :)

OpenStudy (anonymous):

same here :D im an expert at math

OpenStudy (anonymous):

Find the surface area of a small juice can whose diameter is 4 inches and whose height is 6.25 inches. Use 3.14 for π. Hint: The can is a cylinder.

OpenStudy (anonymous):

@mtmb11 help!

OpenStudy (anonymous):

SA=2(PIE)rh+2(PIE)r^2 SA=2(3.14)(2)(6.25)+2(3.14)(2)^2 SA=6.28(12.50)+6.28(4) SA=78.50+25.12 Then just add them together. :)